Nitrogen react with hydrogen to produce ammonia. How many grams Hydrogen is needed, if 10.00 g Nitrogen is used? 
Answer #1

Reaction of nitrogen with hydrogen to form ammonia is as shown

Molar mass of N2=2xmolar mass of N=2x14 g/mol

=28 g/mol

Number of moles of N2 used up=Given mass/molar mass

=10 g/28 g/mol=(10/28) mol

As per the balanced chemical equation given above

1 mol N2 needs 3 mol H2

So (10/28) mol N2 needs 3x(10/28) mol= 1.07 mol H2

Molar mass of H2=2xmolar mass of H=2x1 g/mol=2 g/mol

Mass of hydrogen needed=number of moles x molar mass

=1.07 mol x 2 g/mol=2.14 g

So 2.14 g H2 is needed